OpenStudy (anonymous):

Ms. Soar flew 220 mi in 1.6 h. She then landed at the Aspen Airport to visit a friend. In the second portion of her trip, Ms. Soar flew 312 mi, which took 2.2 h. What was her plane’s average speed for the entire trip?

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

@FireWolfSpirit

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@emmigrace222

OpenStudy (anonymous):

@aleiimmafu

OpenStudy (anonymous):

137.5 + 141.8 = 279.3/2 = 139.65 Answer 139.65 miles per hour ;) @funfuse2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

140

OpenStudy (anonymous):

so is the estimated one correct or the exact?